# A Unified Implementation of Quasi-Monte Carlo Generators, Randomization Routines, and Fast Kernel Methods

### v1

@@ Abstract.
Quasi-random sequences, also called low-discrepancy sequences, have been extensively used as efficient experimental designs across many scientific disciplines. This article provides a unified description and software API for methods pertaining to low-discrepancy point sets. These methods include low discrepancy point set generators, randomization techniques, and fast kernel methods. Specifically, we provide generators for lattices, digital nets, and Halton point sets. Supported randomization techniques include random permutations / shifts, linear matrix scrambling, and nested uniform scrambling. Routines for working with higher-order digital nets and scramblings are also detailed. For kernel methods, we provide implementations of special shift-invariant and digitally-shift invariant kernels along with fast Gram matrix operations facilitated by the bit-reversed FFT, the bit-reversed IFFT, and the FWHT. A new digitally-shift-invariant kernel of higher-order smoothness is also derived. We also describe methods to quickly update the matrix-vector product or linear system solution after doubling the number of points in a lattice or digital net in natural order.
Low-discrepancy sequences, low-discrepancy point set randomizations, higher-order digital net scrambling, randomized Quasi-Monte Carlo, fast kernel methods, shift-invariant kernels, digitally-shift-invariant kernels
copyright: acmlicensedjournalyear: 2025doi: 0.0journal: TOMSjournalvolume: 0journalnumber: 0article: 0publicationmonth: 0submissionid: 0-0-0ccs: Mathematics of computing Statistical softwareccs: Mathematics of computing Mathematical software performanceccs: Mathematics of computing Solversccs: Mathematics of computing Probabilistic algorithms
1. Introduction
This article describes and implements a number of algorithms related to low-discrepancy point sets. These include implementations of
  • low-discrepancy point set generators and randomizations,
  • higher-order digital net generators and randomizations,
  • shift-invariant and digitally-shift-invariant kernels of higher order smoothness,
  • and fast transforms which enable fast Gram matrix operations when special kernels are paired with certain low-discrepancy point sets.
While these methods have been developed and often extensively studied throughout the literature, we found implementations to be scattered or even missing in some popular programming languages. Our main contribution is to provide a unified and accessible software for computations using low-discrepancy sequences in the QMCPy [^1] https://qmcpy.readthedocs.io Python package (Choi et al., 2022b, a). We also integrate with the new LDData [^2] https://github.com/QMCSoftware/LDData repository which provides generating vectors for lattices and generating matrices for digital nets in standardized formats. This collection is, so far, mainly an aggregation of results from the Magic Point Shop [^3] https://people.cs.kuleuven.be/~dirk.nuyens/qmc-generators/ (Kuo and Nuyens, 2016a) and the websites of Frances Kuo on lattices [^4] https://web.maths.unsw.edu.au/~fkuo/lattice/index.html (Cools et al., 2006; Nuyens and Cools, 2006) and Sobol’ points [^5] https://web.maths.unsw.edu.au/~fkuo/sobol/index.html (Joe and Kuo, 2003, 2008).
Below we summarize the novel components of our QMCPy software which we detail in this paper. Comparisons to existing implementations are also discussed. A more comprehensive review of Quasi-Monte Carlo software is given in (Choi et al., 2022b), along with more details on the QMCPy framework.
Lattice point sets:
We support generating points in linear, natural, or Gray code order and randomizing points with shifts modulo one. Other implementations of shifted lattices include those in MATLAB’s GAIL [^6] http://gailgithub.github.io/GAIL_Dev/ (Guaranteed Automatic Integration Library) (Tong et al., 2022; Hickernell et al., 2018), the C++ LatNet Builder [^7] https://github.com/umontreal-simul/latnetbuilder software (L’Ecuyer et al., 2021), and the multi-language MPS [^8] https://people.cs.kuleuven.be/~dirk.nuyens/qmc-generators/ (Magic Point Shop) (Kuo and Nuyens, 2016b). LatNet Builder and MPS also provide search routines for finding good generating vectors which define lattice point sets.
Digital nets in base $2$:
We support generating digital nets in natural or Gray code order and randomizing with linear matrix scrambling (LMS) (Owen, 2003), digital shifts, nested uniform scrambling (NUS) (Owen, 1995), and permutation scrambling. To the best of our knowledge, we are also the only Python implementation to support digital interlacing for the creating of higher order digital nets and higher order scrambling with either LMS or NUS.
@@ 1. Introduction
Early implementations of unrandomized digital sequences, including the Faure, Sobol’ and Niederreiter-Xing constructions, can be found in (Fox, 1986; Bratley et al., 1992; Bratley and Fox, 2003; Pirsic, 2002). Considerations for implementing scrambles for digital sequences was discussed in (Hong and Hickernell, 2003). Recent digital net implementations supporting LMS and digital shifts include those in MPS, PyTorch (Paszke et al., 2019), SciPy (Virtanen et al., 2020), and MATLAB (Inc., 2022). LatNet Builder provides robust support for all randomization routines and digital interlacing for higher order nets and scramblings. LatNet Builder and MPS also provide search routines for finding good generating matrices which define digital nets.
Halton point sets:
@@
As with digital nets, we support randomizing Halton point sets with LMS, digital shifts, NUS, and permutation scrambles. The implementation of Halton point sets and randomizations have been treated in (Owen, 2017; Wang and Hickernell, 2000). The QRNG [^9] https://cran.r-project.org/web/packages/qrng/qrng.pdf (Quasi-Random Number Generators) R package (Hofert and Lemieux, 2023) implements generalized digital nets which use optimized digital permutation scrambles; these are also supported in QMCPy. LatNet Builder also provides support for Halton points and their randomizations.
Kernel Methods and Fast Transforms:
We provide routines to evaluate special shift-invariant and digitally-shift-invariant kernels. This includes a new digitally-shift-invariant kernel of higher-order smoothness derived in this paper. These are not readily available elsewhere.
We also provide routines for fast matrix-vector multiplication and inversion of the nicely-structured Gram matrices produced when pairing these special kernels with certain low-discrepancy sequences. We also describe novel routines to efficiently update the matrix-vector products and linear system solves in the Gram matrix after doubling the number of lattice or digital net points. These fast operations are backed by implementations of the Fast Walsh Hadamard Transform (FWHT), the bit-reversed fast Fourier Transform (FFT), and the bit-reversed inverse FFT (IFFT). We are among the first to provide efficient and accessible Python interfaces to such fast kernel methods and the underlying fast transforms. GAIL has implemented fast Bayesian cubature routines utilizing this machinery.
These above routines are most commonly used in Quasi-Monte Carlo methods and for fast kernel methods, which we describe in the following subsections.
1.1. Quasi-Monte Carlo methods
@@
Monte Carlo (MC) methods approximate a high dimensional integral over the unit cube by the sample average of function evaluations at certain sampling locations:
(1) $\mu:=\int_{[0,1]^{d}}f({\boldsymbol{x}})\mathrm{d}{\boldsymbol{x}}\approx\frac{1}{n}\sum_{i=0}^{n-1}f({\boldsymbol{x}}_{i})=:{\widehat{\mu}}.$
@@
Here $f:[0,1]^{d}\to{\mathbb{R}}$ is a given integrand and $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}\in[0,1]^{n\times d}$ is a point set. The integral on the left hand side may be viewed as taking the expectation ${\mathbb{E}}[f({\boldsymbol{X}})]$ where ${\boldsymbol{X}}$ is a standard uniform ${\boldsymbol{X}}\sim{\mathcal{U}}[0,1]^{d}$. Classic MC methods choose the sampling locations to be independent and identically distributed (IID) $d$-dimensional standard uniforms ${\boldsymbol{x}}_{0},\dots,{\boldsymbol{x}}_{n-1}\overset{\mathrm{IID}}{\sim}{\mathcal{U}}[0,1]^{d}$. IID-MC methods for (1) have a root-mean-squared-error (RMSE) of ${\mathcal{O}}(n^{-1/2})$.
@@
Figure 1. An independent identically distributed (IID) point set and low-discrepancy (quasi-random) point sets of size $n=2^{13}=8192$. Notice the quasi-random points more evenly fill the space than IID points, which leads to faster convergence of Quasi-Monte Carlo methods compared to IID Monte Carlo methods. Randomized lattices, digital nets in base $b=2$ (DN2), and Halton points are shown. Randomizations include shifts, linear matrix scramblings (LMS), digital shifts (DS), and nested uniform scramblings (NUS). Digital interlacing of order $\alpha$ is used to create higher-order randomized digital nets (HOα).
[…]
@@
Quasi-Monte Carlo (QMC) methods (Niederreiter, 1992; Dick and Pillichshammer, 2010; Kroese et al., 2013; Dick et al., 2022; Lemieux, 2009; Sloan and Joe, 1994) replace IID point sets with low-discrepancy (LD) point sets which more evenly cover the unit cube $[0,1]^{d}$. For integrands with bounded variation, plugging LD point sets into ${\widehat{\mu}}$ (1) yields a worst-case error rate of ${\mathcal{O}}(n^{-1+\delta})$ with $\delta>0$ arbitrarily small. Some popular LD point sets are plotted in fig. 1 including lattices, digital nets (including higher-order versions), and Halton point sets.
@@
Randomized Quasi-Monte Carlo (RQMC) uses randomized LD point sets to give improved convergence rates and enable practical error estimation. Specifically, if we again assume the integrand has bounded variation, then certain RQMC methods can achieve a RMSE of ${\mathcal{O}}(n^{-3/2+\delta})$. Moreover, let $\{{\boldsymbol{x}}_{i}^{1}\}_{i=0}^{n-1},\dots,\{{\boldsymbol{x}}_{i}^{R}\}_{i=0}^{n-1}$ denote $R$ IID randomizations of a low-discrepancy point set where typically $R$ is small e.g. $R=15$. Then the RQMC estimate
(2) ${\widehat{\mu}}_{R}=\frac{1}{R}\sum_{r=1}^{R}\hat{\mu}^{r}\qquad\text{where}\quad\hat{\mu}^{r}=\frac{1}{n}\sum_{i=0}^{n-1}f({\boldsymbol{x}}_{i}^{r})$
may be used to construct a practical (approximate) $100(1-\tau)\%$ confidence interval ${\widehat{\mu}}_{R}\pm t_{R-1,\tau/2}{\widehat{\sigma}}_{R}/\sqrt{R}$ for $\mu$ where ${\widehat{\sigma}}_{R}^{2}=R^{-1}(R-1)^{-1}\sum_{r=1}^{R}({\widehat{\mu}}^{r}-{\widehat{\mu}}_{R})^{2}$ and $t_{R-1,\tau/2}$ is the $\tau/2$ quantile of a Student distribution with $R-1$ degrees of freedom. RQMC estimates are detailed in (Owen, 2018, Chapter 17), and a recent comparison of RQMC confidence interval methods can be found in (L’Ecuyer et al., 2023). QMC error estimation is treated more broadly in (Owen, 2024), while (Clancy et al., 2014; Hickernell et al., 2017; Clancy et al., 2014) detail additional considerations for adaptive QMC algorithms. For lattices, randomization is typically done using random shifts modulo $1$. For digital nets, randomization is typically done using nested uniform scrambling (NUS) or the cheaper combination of linear matrix scrambling (LMS) with digital shifts / permutations (Owen, 2003, 2017; Owen and Pan, 2024).
@@
Higher-order LD point sets were designed to yield faster convergence for integrands with additional smoothness. For integrands with square integrable mixed partial derivatives up to order $\alpha>1$, plugging higher-order digital nets into ${\widehat{\mu}}$ (1) yields a worst-case error rate of ${\mathcal{O}}(n^{-\alpha+\delta})$ (Dick, 2008, 2009b, 2009a). RQMC using scrambled digital nets (via either NUS or LMS with digital permutations) combined with digital interlacing has been shown to achieve an RMSE of order ${\mathcal{O}}(n^{-\alpha-1/2+\delta})$ (Dick, 2011).
1.2. Kernel computations

### v2

# QMCPy: A Python Software for Randomized Low-Discrepancy Sequences, Quasi-Monte Carlo, and Fast Kernel Methods

@@ Abstract
Low-discrepancy (LD) sequences have been extensively used as efficient experimental designs across many scientific disciplines. QMCPy (https://qmcsoftware.github.io/QMCSoftware/) is an accessible Python library which provides a unified implementation of randomized LD sequences, automatic variable transformations, adaptive Quasi-Monte Carlo error estimation algorithms, and fast kernel methods. This article focuses on recent updates to QMCPy which broaden support for randomized LD sequences and add new tools to enable fast kernel methods using LD sequences. Specifically, we give a unified description of the supported LD lattices, digital nets, and Halton point sets, along with randomization options including random permutations / shifts, linear matrix scrambling (LMS), and nested uniform scrambling (NUS). We also support higher-order digital nets, higher-order scrambling with LMS or NUS, and Halton scrambling with LMS or NUS. For fast kernel methods, we provide shift-invariant (SI) and digitally-shift-invariant (DSI) kernels, including a new set of higher-order smoothness DSI kernels. When SI and DSI kernels are respectively paired with $n$ LD lattice and digital net points, the resulting Gram matrices permit multiplication and inversion at only $\mathcal{O}(n\log n)$ cost. These fast operations utilize QMCPy’s implementation of the fast Fourier transform in bit-reversed order (FFTBR), inverse FFTBR (IFFTBR), and fast Walsh–Hadamard transform (FWHT).
1 Introduction
Low-discrepancy (LD) sequences, also called quasi-random sequences, judiciously explore the unit cube in high dimensions. They were first developed to replace the independent points used in Monte Carlo simulation for high dimensional numerical integration. The resulting Quasi-Monte Carlo (QMC) methods have proven theoretically superior to independent Monte Carlo methods for a large class of nicely behaved functions [Niederreiter, 1992, Dick and Pillichshammer, 2010, Kroese et al., 2013, Dick et al., 2022, Lemieux, 2009, Sloan and Joe, 1994, Dick et al., 2013]. This advantage has been shown across a variety of scientific disciplines including financial modeling [Joy et al., 1996, Lai and Spanier, 1998, L’Ecuyer, 2004, 2009, Xu and Ökten, 2015, Giles and Waterhouse, 2009], solving PDEs with random coefficients [Graham et al., 2011, Kuo et al., 2012, 2015, Graham et al., 2015, Kuo and Nuyens, 2016, Robbe et al., 2017], and graphics rendering using ray tracing [Jensen et al., 2003, Raab et al., 2006, Waechter and Keller, 2011] among others. Randomizing LD sequences provide additional advantages, such as enabling practical QMC error estimation, avoiding boundary observations, and negating the probability of pairing to adverse functions [Owen, 1995, L’Ecuyer, 2016, L’Ecuyer et al., 2023, Owen, 2003, Matoušek, 1998, Tezuka, 2002].
LD sequences, especially randomized versions, have also been extensively used as efficient experimental designs. In reproducing kernel Hilbert spaces (RKHSs), pairing certain LD points with certain kernels yields structured Gram matrices which are diagonalizable by fast transforms [Zeng et al., 2009, 2006]. For example, pairing an LD rank-1 lattice in linear order with a shift-invariant kernel gives a circulant Gram matrix which is diagonalizable by the fast Fourier transform (FFT) and inverse FFT (IFFT). Such structures have been exploited to accelerate kernel interpolation methods for solving PDEs with random coefficients [Kaarnioja et al., 2022, 2023, Sorokin et al., 2024], Bayesian cubature [Rathinavel and Hickernell, 2019, 2022, Rathinavel, 2019], and discrepancy measure computations [Hickernell, 1998a, b]. LD points have also found recent applications to training scientific machine learning models [Longo et al., 2021, Chen et al., 2021, Keller et al., 2025], although this is an avenue we do not discuss in this paper.
While randomized LD sequences have been extensively studied for QMC and fast kernel methods, implementations of such methods are often scattered or even missing in some popular programming languages. Our prior works [Choi et al., 2022b, a, Hickernell et al., 2025, Sorokin and Rathinavel, 2022] described the current QMC software landscape and introduced QMCPy: a community driven implementation of QMC routines into a unified and accessible Python framework.
1.1 QMCPy features
Below we summarize the supported features in QMCPy [^1] https://qmcsoftware.github.io/QMCSoftware/ and provide pointers to related codes:
Randomized LD Sequences
We support the point sets and randomization routines described below. These features are also supported in the comprehensive C++ LatNet Builder [^2] https://github.com/umontreal-simul/latnetbuilder software [L’Ecuyer et al., 2021]. LatNet Builder and the multi-language Magic Point Shop (MPS) [^3] https://people.cs.kuleuven.be/~dirk.nuyens/qmc-generators/ [Kuo and Nuyens, 2016] both provide search routines for finding good lattice generating vectors and digital net generating matrices. QMCPy integrates with the new LDData [^4] https://github.com/QMCSoftware/LDData repository which contains a variety of these pregenerated vectors and matrices in standardized formats. LDData additionally includes popular choices from the websites of Frances Kuo on lattices [^5] https://web.maths.unsw.edu.au/~fkuo/lattice/index.html [Cools et al., 2006, Nuyens and Cools, 2006] and Sobol’ points [^6] https://web.maths.unsw.edu.au/~fkuo/sobol/index.html (a special case of digital nets) [Joe and Kuo, 2003, 2008].
Lattice Points
Rank-1 lattices may be generated in radical inverse or linear order, and may be randomized using shifts modulo one. Randomly shifted rank-1 lattices are also available in MPS and GAIL [^7] http://gailgithub.github.io/GAIL_Dev/ (MATLAB’s Guaranteed Automatic Integration Library) [Tong et al., 2022, Hickernell et al., 2018].
Digital Nets
@@ 1.1 QMCPy features
Base $2$ digital nets, including higher-order digital nets, may be generated in either radical inverse or Gray code order, and may be randomized with linear matrix scrambling (LMS) [Owen, 2003], digital shifts, nested uniform scrambling (NUS) [Owen, 1995], and/or permutation scrambling. Early implementations of unrandomized digital sequences, including the Faure, Sobol’, and Niederreiter–Xing constructions, can be found in [Fox, 1986, Bratley et al., 1992, Bratley and Fox, 2003, Pirsic, 2002]. Considerations for implementing scrambles were discussed in [Hong and Hickernell, 2003]. Support for combining LMS with digital shifts is also provided in MATLAB, MPS, and both the PyTorch [^8] https://pytorch.org/ [Paszke et al., 2019] and SciPy [^9] https://scipy.org/ [Virtanen et al., 2020] Python packages.
Halton Points
@@
As with digital nets, Halton point sets may be randomized with LMS, digital shifts, NUS, and/or permutation scrambles. The implementation of Halton point sets and randomizations have been treated in [Owen, 2017, Wang and Hickernell, 2000]. The QRNG [^10] https://cran.r-project.org/web/packages/qrng/qrng.pdf (Quasi-Random Number Generators) R package [Hofert and Lemieux, 2023] implements generalized Halton point sets [Faure and Lemieux, 2009] which use optimized digital permutation scrambles; these are also supported in QMCPy.
Variable Transformations
These define the distribution of stochasticity in the underlying problem and automatically rewrite user-defined functions into QMC-compatible forms. The available transforms are mainly wrappers around distributions provided by the SciPy Python package.
Error Estimators for (Q)MC and Multilevel (Q)MC
We provide numerous adaptive error estimation algorithms which automatically select the number of points required for a (Q)MC approximation to be within user-specified error tolerances. QMC error estimation is treated more broadly in [Owen, 2025], while [Clancy et al., 2014, Hickernell et al., 2017] detail additional considerations for adaptive QMC algorithms. QMCPy’s adaptive error estimation procedures are described below. The single level algorithms were also implemented in GAIL.
Monte Carlo with Independent Points
Confidence intervals are derived from either a central limit theorem (CLT) heuristic or a guaranteed version of CLT for functions with bounded kurtosis [Hickernell et al., 2013].
QMC with Multiple Randomizations
Student’s-$t$ confidence intervals are evaluated based on independent mean estimates from independent randomizations of an LD point set, see [L’Ecuyer et al., 2023] or [Owen, 2023, Chapter 17].
QMC via Decay Tracking using a Single Randomized LD Sequence
Guaranteed error bounds are available for cones of functions whose Fourier or Walsh coefficients decay predictably [Hickernell and Jiménez Rugama, 2014, Hickernell et al., 2017, Jiménez Rugama and Hickernell, 2014, Ding et al., 2018].
QMC via Bayesian Cubature using a Single Randomized LD Sequence
Posterior credible intervals on the integral of a Gaussian process may be computed quickly by exploiting fast kernel computations [Rathinavel and Hickernell, 2019, 2022, Rathinavel, 2019].
Multilevel Monte Carlo with Independent Points
We have implemented standard multilevel Monte Carlo [Giles, 2008] and continuation multilevel Monte Carlo algorithms [Collier et al., 2015, Robbe et al., 2016, 2019].
Multilevel QMC with Multiple Randomizations
We have implemented standard multilevel QMC [Giles and Waterhouse, 2009] and continuation multilevel QMC algorithms [Robbe et al., 2017].
Fast Kernel Computations
Matching a rank-1 lattice in radical inverse order to a shift-invariant (SI) RKHS kernel yields a Gram matrix diagonalizable by the FFT in bit-reversed order (FFTBR) and its inverse (IFFTBR). Similarly, matching a base $2$ digital net in radical inverse order to a digitally-shift-invariant (DSI) kernel yields a Gram matrix diagonalizable by the fast Walsh–Hadamard transform (FWHT) [Fino and Algazi, 1976]. The currently supported kernels and fast transforms are described below.
RKHS Kernels
We provide SI kernels and DSI kernels, including those of higher-order smoothness. We also maintain an interface to a number of standard kernels, such as squared exponential and Matérn kernels, whose implementations may also be found in the GPyTorch [^11] https://gpytorch.ai/ [Gardner et al., 2018] and the scikit-learn [^12] https://scikit-learn.org/ [Pedregosa et al., 2011] Python packages among others. SI kernels of arbitrary smoothness are well known and can be computed based on the Bernoulli polynomials [Kaarnioja et al., 2022, 2023, Cools et al., 2021, 2020, Sloan and Woźniakowski, 2001, Kuo et al., 2004]. DSI kernels of order $1$ smoothness were derived in [Dick and Pillichshammer, 2005]. In this paper we derive new higher-order DSI kernels whose RKHSs contain smooth functions.
Fast Transforms
We provide interfaces to the FFTBR, IFFTBR, and FWHT algorithms which all have $\mathcal{O}(n\log n)$ complexity in the number of points $n$. Our FFTBR and IFFT implementations use the FFT routines in the SciPy. An implementation of the FWHT is also available in the SymPy [^13] https://www.sympy.org [Meurer et al., 2017] Python package.
1.2 Novel contributions and scope
The novel contributions of this article are to:
  1. 1.
    Update QMCPy to be the first Python interface to support higher-order digital nets, higher-order digital net scrambling with either LMS or NUS, or Halton scrambling with either LMS and NUS.
  2. 2.
    Introduce the LDData repository, which contains collections of generating vectors and generating matrices in standardized formats, and provide a supported interface in QMCPy.
  3. 3.
    Add new tools to QMCPy which enable fast kernel computations. These tools include SI and DSI kernels of varying smoothness as well as the FFTBR, IFFTBR, and FWHT algorithms.
  4. 4.
    Derive novel DSI kernels of higher-order smoothness. Low order kernel forms appeared in [Baldeaux et al., 2012] as worst-case error bounds on QMC rules using higher-order polynomial lattices, but there they were not interpreted as DSI kernels. We also derive a new order $4$ smoothness DSI kernel whose form has not appeared elsewhere in the literature.
  5. 5.
    Provide efficient procedures to update the FFTBR, IFFTBR, or FWHT of a sequence when the number of points is doubled and the original points are not discarded.
This article will focus on our novel contributions to equip QMCPy with broad support for randomized LD sequences and new tools for fast kernel methods.
1.3 Outline
The remainder of this article is organized as follows. Section 2 gives common notations. Section 3 describes two motivating problems using low discrepancy sequences: QMC methods (Section 3.1) and fast kernel computations (Section 3.2). Section 4 gives a few notes on our QMCPy implementation and the setup required to reproduce codes in this article. Section 5 details randomized rank-1 lattices (Section 5.1), digital sequences (Section 5.2), and Halton sequences (Section 5.3). Section 6 describes pairing SI kernels to rank-1 lattices (Section 6.1), DSI kernels to digital nets (Section 6.2), and the resulting fast kernel computations (Section 6.3). Section 7 provides numerical experiments showcasing the speed, accuracy, and versatility of QMCPy. Section 8 gives a brief conclusion.
2 Notation
Bold symbols will denote vectors which are assumed to be column vectors, e.g., ${\boldsymbol{x}}\in[0,1]^{d}$. Capital letters in serif font denote matrices, e.g., ${\mathsf{K}}\in{\mathbb{R}}^{n\times n}$. Lower case letters in serif font denote digits in a base $b$ expansion, e.g., $i=\sum_{t=0}^{m-1}{\mathsf{i}}_{t}b^{t}$ for $0\leq i<2^{m}$. This may be combined with bold notation when denoting the vector of base $b$ digits, e.g., ${\boldsymbol{\mathsf{i}}}=({\mathsf{i}}_{0},{\mathsf{i}}_{1},\dots,{\mathsf{i}}_{m-1})^{\intercal}$. Modulo is always taken component-wise, e.g., rank-1 lattices will use the notation ${\boldsymbol{x}}\mod 1=(x_{1}\mod 1,\dots,x_{d}\mod 1)^{\intercal}$ and digital nets will take all matrix operations to be carried modulo $b$. Permutations may be specified by vectors, e.g., $\pi:\{0,1,2\}\to\{0,1,2\}$ with $\pi(0)=2$, $\pi(1)=0$, and $\pi(2)=1$ may be denoted by $\pi=(2,0,1)$.
3 Motivating problems
3.1 QMC methods
@@
Monte Carlo (MC) and Quasi-Monte Carlo (QMC) methods approximate a high dimensional integral over the unit cube by the sample average of function evaluations at certain sampling locations:
$\mu:=\int_{[0,1]^{d}}f({\boldsymbol{x}})\mathrm{d}{\boldsymbol{x}}\approx\frac{1}{n}\sum_{i=0}^{n-1}f({\boldsymbol{x}}_{i})=:{\widehat{\mu}}$. (1)
@@
Here $f:[0,1]^{d}\to{\mathbb{R}}$ is a given integrand and $\{{\boldsymbol{x}}_{i}\}_{i=0}^{n-1}\in[0,1]^{n\times d}$ is a point set. The integral on the left-hand side may be viewed as taking the expectation ${\mathbb{E}}[f({\boldsymbol{X}})]$ where $\boldsymbol{X}$} is a standard uniform ${\boldsymbol{X}}\sim{\mathcal{U}}[0,1]^{d}$. Classic MC methods choose the sampling locations to be independent and identically distributed (IID) $d$-dimensional standard uniforms ${\boldsymbol{x}}_{0},\dots,{\boldsymbol{x}}_{n-1}\overset{\mathrm{IID}}{\sim}{\mathcal{U}}[0,1]^{d}$. IID MC methods for (1) have a root-mean-squared-error (RMSE) of ${\mathcal{O}}(n^{-1/2})$.
@@
Figure 1: An independent identically distributed (IID) point set and low-discrepancy (LD) point sets of size $n=2^{13}=8192$. Notice the LD points more evenly fill the space than IID points, which leads to faster convergence of Quasi-Monte Carlo methods compared to IID Monte Carlo methods. Randomized rank-1 lattices, digital nets in base $2$, and Halton points are shown. Randomizations include shifts, linear matrix scramblings (LMS), digital shifts (DS), digital permutations (DP), and nested uniform scramblings (NUS). Digital interlacing of order $\alpha$ is used to create higher-order randomized digital nets in base 22 (DNα).
@@
QMC methods [Niederreiter, 1992, Dick and Pillichshammer, 2010, Kroese et al., 2013, Dick et al., 2022, Lemieux, 2009, Sloan and Joe, 1994, Dick et al., 2013] replace IID point sets with LD point sets which more evenly cover the unit cube $0,1]^{d}$. For integrands with bounded variation, plugging LD point sets into (1) yields a worst-case error rate of ${\mathcal{O}}(n^{-1+\delta})$ with $\delta>0$ arbitrarily small. Some popular LD point sets are plotted in Figure 1 including randomized rank-1 lattices, base $2$ digital nets (including higher-order versions), and Halton point sets.
@@
Randomized Quasi-Monte Carlo (RQMC) uses randomized LD point sets to give improved convergence rates and enable practical error estimation. Specifically, if we again assume the integrand has bounded variation, then certain RQMC methods can achieve a RMSE of ${\mathcal{O}}(n^{-3/2+\delta})$. Moreover, let $\{{\boldsymbol{x}}_{1i}\}_{i=0}^{n-1},\dots,\{{\boldsymbol{x}}_{Ri}\}_{i=0}^{n-1}$ denote $R$ IID randomizations of an LD point set where typically $R$ is small, e.g., $R=15$. Then, following [L’Ecuyer et al., 2023] and [Owen, 2023, Chapter 17], the RQMC estimate
${\widehat{\mu}}=\frac{1}{R}\sum_{r=1}^{R}{\widehat{\mu}}_{r}\qquad\text{where}\quad{\widehat{\mu}}_{r}=\frac{1}{n}\sum_{i=0}^{n-1}f({\boldsymbol{x}}_{ri})$ (2)
gives a practical $100(1-\tau)\%$ confidence interval ${\widehat{\mu}}\pm t_{R-1,\tau/2}{\widehat{\sigma}}/\sqrt{R}$ for $\mu$ where ${\widehat{\sigma}}^{2}=(R-1)^{-1}\sum_{r=1}^{R}({\widehat{\mu}}_{r}-{\widehat{\mu}})^{2}$ and $t_{R-1,\tau/2}$ is the $\tau/2$ quantile of a Student’s-$t$ distribution with R1R-1 degrees of freedom. LABEL:code:genz_ex_1 will show how to code up such an estimate in QMCPy. For rank-1 lattices, randomization is typically done using random shifts modulo $1$. For digital nets, randomization is typically done using NUS or the cheaper combination of LMS with digital shifts / permutations [Matoušek, 1998, Owen, 2003, 2017, Owen and Pan, 2024].
@@
Higher-order LD point sets were designed to yield faster convergence for integrands with additional smoothness. For integrands with square integrable mixed partial derivatives up to order $\alpha>1$, plugging higher-order digital nets into $^{\widehat{\mu}}$ (1) yields a worst-case error rate of ${\mathcal{O}}(n^{-\alpha+\delta})$ [Dick, 2008, 2009b, 2009a]. RQMC using higher-order digital nets with higher-order NUS or LMS scrambling has been shown to achieve an RMSE of order ${\mathcal{O}}(n^{-\alpha-1/2+\delta})$ [Dick, 2011].
3.2 Fast Kernel computations